A rectangular building contains 3600 square ft. How many feet are in the perimeter of the lot?
Answer by checkley77(12844)

This problem has an infinite number of solutions.
Im assuming the building is the same as the lot size.
3,600=xy
Let x=60 & y=60
60*60=3,600
Then the perimeter would be:
2*60+2*60=120+120=240 ft.
------------------------------
But let x=120 & y=30
120*30=3,600
Now the perimeter is:
2*120+2*30=240+50=290 ft.
-------------------------------
But let x=100 & y=36
100*36=3,6000
Now the perimeter is:
2*100+2*36=200+72=272 ft.